OpenStudy (anonymous):

For the function f(x) = (3 – 4x)², find f–1. Determine whether f–1 is a function. show work. help. medal rewaarded!

OpenStudy (anonymous):

f(-1) ? or (f-1) Which do you mean

OpenStudy (anonymous):

its f^-1

OpenStudy (anonymous):

f(x)^-1?

OpenStudy (anonymous):

If so then it would just be f(x)^-1=(3-4x) because the exponents when multiplying are added. 2+(-1)=1 Unless I have interpreted what the question is asking for wrongly.

OpenStudy (anonymous):

No... Switch x and y then solve for y. y = (3 – 4x)² x = (3-4y)^2 Solve for y

OpenStudy (anonymous):

Oh right that means inverse lol. Yes, make sure it's inversable as a one to one function, switch x and y like ^^^^said and solve for y.

OpenStudy (anonymous):

im stumped!:(

OpenStudy (anonymous):

x = (3-4y)^2 y= 1/4 (3-sqrt(x)) so... f^-1 = 1/4 (3-sqrt(x))

OpenStudy (anonymous):

@kudrick16